### v1

# Grounded Reinforcement Learning for Visual Reasoning

Abstract
While reinforcement learning (RL) over chains of thought has significantly advanced language models in tasks such as mathematics and coding, visual reasoning introduces added complexity by requiring models to direct visual attention, interpret perceptual inputs, and ground abstract reasoning in spatial evidence. We introduce ViGoRL (Visually Grounded Reinforcement Learning), a vision-language model trained with RL to explicitly anchor each reasoning step to specific visual coordinates. Inspired by human visual decision-making, ViGoRL learns to produce spatially grounded reasoning traces, guiding visual attention to task-relevant regions at each step. When fine-grained exploration is required, our novel multi-turn RL framework enables the model to dynamically zoom into predicted coordinates as reasoning unfolds. Across a diverse set of visual reasoning benchmarks—including SAT-2 and BLINK for spatial reasoning, Vbench for visual search, and ScreenSpot and VisualWebArena for web-based grounding—ViGoRL consistently outperforms both supervised fine-tuning and conventional RL baselines that lack explicit grounding mechanisms. Incorporating multi-turn RL with zoomed-in visual feedback significantly improves ViGoRL’s performance on localizing small GUI elements and visual search, achieving 86.4% on VBench. Additionally, we find that grounding amplifies other visual behaviors such as region exploration, grounded subgoal setting, and visual verification. Finally, human evaluations show that the model’s visual references are not only spatially accurate but also helpful for understanding model reasoning steps. Our results show that visually grounded RL is a strong paradigm for imbuing models with general-purpose visual reasoning.
Figure 1: Grounded visual reasoning enables interpretable and accurate answers. ViGoRL decomposes the task into a sequence of natural language thoughts anchored in image regions. In contrast, Vanilla GRPO and SFT baselines produce ungrounded and incorrect responses.
1 Introduction
Visual reasoning tasks vary widely in structure and often demand different solution strategies depending on the problem at hand. Some tasks are dominated by salient visual cues, such as recognizing a refrigerator centered in a kitchen scene, while others, like locating a pair of scissors in a cluttered environment, require sequential visual search and selective attention. Despite this diversity, most state-of-the-art vision-language models (VLMs) operate in an end-to-end fashion, predicting answers directly in a single forward pass. These models often lack the ability to adapt their computational strategies to different tasks or to expose intermediate reasoning beyond visual attention maps. Prompt-based models, such as ViperGPT [54], VisualProg [19], and V* [67], explicitly decompose visual tasks into sequences of subgoals or intermediate steps to improve interpretability and performance without additional training. However, such models typically generate fixed reasoning chains that do not adapt to the structure of the input scene.
Recent advances in reinforcement learning (RL) over reasoning chains have significantly enhanced the capabilities of LLMs in text-based domains [17, 58, 26], enabling them to learn diverse reasoning strategies tailored to the context. However, RL can only build upon skills or compose reasoning behaviors that are already latent in the base model’s sampling distribution [13, 82]. For e.g., Gandhi et al. [13] has identified key cognitive behaviors in text-based domains, such as setting sub-goals, backtracking, verification, that support self-improvement under RL. Models lacking these behaviors often do not benefit from RL and must be bootstrapped via supervised fine-tuning (SFT) on curated reasoning traces before RL is run [13, 82]. However, it remains unclear whether the cognitive behaviors identified in text-based domains similarly support generalization in visual reasoning tasks.
Several recent works have attempted RL fine-tuning directly on base vision-language models (VLMs) [38, 37, 87, 55, 57, 84, 51, 41, 32], implicitly assuming RL alone can induce useful cognitive behaviors. However, our analysis reveals that such naïve applications of RL typically yield abstract, ungrounded reasoning rather than richer, visually grounded cognitive behaviors (see Section 3.1,  5.3). These findings align with prior research showing that explicitly prompting VLMs to reference spatial object locations improves performance and interpretability [67, 73, 14], suggesting that grounding thought in spatial regions may serve as a key cognitive behavior for effective visual reasoning. Thus, a critical open question arises: How can we embed useful cognitive behaviors in VLMs before applying RL to achieve robust visual reasoning?
We hypothesize that models both “see better” and “think better” when their textual reasoning steps are explicitly grounded in specific image regions, promoting more targeted and systematic cross-referencing between textual and visual information during reasoning. This hypothesis is inspired by the fact that humans systematically shift their restricted gaze to selectively gather and integrate task-relevant information when reasoning about the world [75, 22, 74]. Grounding may serve a similar role in models, functioning as a spatial attention mechanism that enables accurate feature binding [61] and supports deictic reference [4] to simplify multi-step reasoning through localized perceptual anchoring. We move beyond prompt-based reasoning, proposing that learning to compose reasoning steps explicitly anchored in image coordinates induces structured region-level behaviors that support improved generalization in visual tasks.
Our Approach. We introduce a multi-turn RL framework for training VLMs to reason in a grounded, visually-aware manner. This stands in contrast to LLM reasoning in math or code, where grounding in external input is not strictly required. Within each reasoning step, the model produces a natural language thought along with a corresponding spatial grounding (i.e., an $(x,y)$ location in the image). This enables it to progressively refine its attention and gather task-relevant visual information as reasoning unfolds. By incorporating multi-turn interaction into the RL process—where each turn consists of one or more reasoning steps followed by a query to a visual feedback tool—the model learns to iteratively request zoomed-in views of selected regions when fine-grained visual information is required. Critically, no external supervision or explicit human-provided grounding cues are used to supervise the spatial grounding of the thought; instead, the model autonomously learns to propose and utilize spatial grounding as an internal cognitive tool.
Current methods for training VLMs to directly produce textual answers from visual inputs inherently bias them toward abstract, ungrounded reasoning, making it fundamentally difficult for RL methods alone to spontaneously discover systematic visual strategies at the region-level. To explicitly inject grounded reasoning behaviors before RL training, we employ Monte Carlo Tree Search (MCTS) to systematically stitch together independently sampled reasoning steps, generating diverse, visually-grounded reasoning trajectories. We bootstrap the model via supervised fine-tuning (SFT) on these MCTS-constructed paths, thus embedding rich region-level reasoning strategies into the model. We then apply RL, through Group Relative Policy Optimization (GRPO) [50], to further reinforce grounded sequences that lead to correct answers. Finally, we introduce a novel multi-turn RL formulation with visual feedback loops, allowing the model to dynamically zoom into image regions via tool calling for more detailed visual inspection when needed. This multi-turn variant of our method improves the model’s capacity to localize and reason about fine-grained visual elements.
Empirical Results. We evaluate ViGoRL across a suite of visual reasoning benchmarks, including SAT-2 [48], BLINK [11], RoboSpatial [53], ScreenSpot [6, 31], VisualWebArena [29], and VBench [67]. Our approach consistently outperforms existing methods on all tasks. Specifically, ViGoRL achieves substantial improvements over vanilla GRPO, with accuracy gains of 12.9 points on SAT-2 and 2.0 points on BLINK. In fine-grained web grounding scenarios, our method surpasses both vanilla GRPO and large-scale web-finetuned models on ScreenSpot-Pro. By leveraging multi-turn RL for dynamic, zoomed-in visual feedback, ViGoRL further improves performance on ScreenSpot-Pro, effectively localizing small elements within high-resolution images. Moreover, multi-turn RL significantly enhances visual search capabilities, allowing ViGoRL to outperform both VLM tool-use pipelines and proprietary VLMs on VBench, achieving an accuracy of 86.4%. On VisualWebArena, a benchmark requiring live web interaction from image inputs alone, without access to HTML, ViGoRL outperforms both direct SFT and vanilla GRPO, and surpasses the previous state-of-the-art for this model size, ICAL [49], despite using only visual input.
Ablation studies confirm the importance of grounding: models trained without spatial anchoring perform significantly worse. Further, we find that grounding amplifies other visual cognitive behaviors such as region exploration, goal setting, and visual verification. Human evaluations show that our model’s reasoning is both spatially accurate and helpful to understanding the model’s reasoning steps. Our results point to visually grounded RL as a strong paradigm for general-purpose visual reasoning.
2 Related Work
Programmatic Reasoning in VLMs. Vision-Language Models (VLMs) [46, 27, 2, 5, 63, 56] excel on multimodal tasks through large-scale pretraining, but struggle with complex reasoning such as counting [47], spatial reasoning [48], and compositional understanding [60]. Prompting strategies like chain-of-thought (CoT) [65, 30], mm-CoT [86], IoT prompting [91], and Mind’s Eye [69] guide models to generate explicit reasoning steps grounded in images. Methods such as V* [67], Sketchpad [24], VisProg [19], REFOCUS [12], and ViperGPT [54] use language models to produce executable visual plans but rely on frozen backbones and hand-crafted prompts.
Distillation and Supervised Fine-Tuning. To robustly instill reasoning skills, supervised fine-tuning (SFT) methods train on curated reasoning trajectories. For text tasks, STaR [83] generates reasoning via few-shot prompting and selects based on correctness, while S1 [40] distills reasoning chains into smaller models. Similar approaches in VLMs include LLAVA-CoT [72], which distills CoT reasoning from GPT-4o [1], and ICAL [49], VPD [25], and Mulberry [77], which employ LLMs or MCTS to generate reasoning data. Methods like VOCOT [33] improve grounding of entities via SFT. However, distillation methods rely solely on positive examples, neglecting failed paths. RL addresses this by learning from both successes and failures, outperforming SFT alone in our experiments.
Reinforcement Learning on Chains of Thought. Applying RL to chains of thought has improved reasoning in verifiable domains like math and coding [26, 17, 58]. Early methods [85, 43, 10, 68, 16] iteratively refine reasoning using preference-based approaches, while recent efforts like DeepSeek-R1 [17] and Kimi-1.5 [58] leverage online RL with outcome-based rewards. While initially believed to induce novel cognitive behaviors, analyses [13, 82, 36, 79] suggest RL primarily amplifies existing capabilities already found in the base model. Most prior work focuses on text-only domains, leaving visual reasoning behaviors largely unexplored. Visual-RFT [37] applies RL to textual reasoning in VLMs without incentivizing new visual behaviors. In contrast, we explicitly ground reasoning steps visually, amplifying exploration, verification, and backtracking via learned visual interactions.
3 Preliminaries
Reasoning in Vision-Language Models. The overarching objective of our work is to improve the reasoning capabilities of vision-language models (VLMs). We consider reasoning tasks defined by a dataset $\mathcal{D}$ of problem instances $(I,q,a^{\ast})$, where $I$ is a visual input (e.g., an image), $q$ is a natural language query about the image, and $a^{\ast}$ is the correct, verifiable answer (e.g., a class label, bounding box, or discrete actions such as click [element]). The goal is to train a vision-language policy $\pi_{\theta}$ parameterized by $\theta$ that outputs a reasoning trace $\tau$ consisting of sequential textual reasoning steps $s_{1},s_{2},\dots,s_{T}$ culminating in an answer $a$. This policy factorizes autoregressively:
$\pi_{\theta}\bigl{(}\tau\mid I,q\bigr{)}\;=\;\Bigl{(}\prod_{t=1}^{T}\pi_{\theta}\bigl{(}s_{t}\mid I,q,s_{<t}\bigr{)}\Bigr{)}\;\cdot\;\pi_{\theta}\bigl{(}a\mid I,q,s_{\leq T}\bigr{)}.$
While supervised fine-tuning can teach models to mimic reasoning chains provided in training data, RL offers the potential to directly reinforce reasoning behavior sampled from the base model based on correctness or other reward signals [42, 50, 13]. Specifically, RL allows us to optimize policies over reasoning traces $\tau$ that maximize expected returns based on task performance and adherence to desired format structure. Formally, the RL objective can be expressed as:
$\max_{\theta}\;\mathbb{E}_{(I,q,a^{\ast})\sim\mathcal{D}}\;\left[\mathbb{E}_{\tau\sim\pi_{\theta}}\bigl{[}\,R(\tau)\bigr{]}\right],\quad\text{s.t.}\;\tau\text{ satisfies format constraints},$
where the reward $R(\tau)$ typically includes correctness of the final answer, and proper adherence to structured reasoning formats.
3.1 Do Current RL Recipes Amplify VLM Behaviors That Support Visual Performance?
It has been shown that RL on chains of thought alone cannot necessarily induce new behaviors from scratch; it can only amplify or chain reasoning primitives that are already present in the base model’s sampling distribution [13, 82]. Do current base VLMs exhibit desirable visual reasoning behaviors and can RL amplify these behaviors to improve performance? Following previous work from LLMs [13], we categorize the behaviors of base VLMs when tasked with the Spatial Aptitude Test [48] spatial reasoning benchmark, which requires the synthesis of evidence in multiple regions of one or multiple images to answer. Our complete analysis and experimental protocol is detailed in Section 5.3 and Appendix A6. Our analysis reveals two key insights:
  • Takeaway 1: Current VLMs often fail to reference fine-grained image inputs; their reasoning is largely ungrounded. Without explicit grounding, models treat vision as static context rather than actively referenced input, spending tokens on abstract thought instead of analyzing visual regions. Qwen2.5-VL-3B examines only 1.44 regions per task with minimal visual verification (0.14) and no backtracking. We provide an examples of such an output in Figure 2.
  • Takeaway 2: Standard RL optimization exacerbates ungrounded reasoning. RL-tuning with task-level rewards slightly increases region exploration (1.8) but eliminates visual subgoal setting (0.00) and still shows no backtracking (0.00). Figure 2 and Appendix A9.1 illustrates typical outputs under this setup. This demonstrates that optimizing for correctness without encouraging grounded reasoning fails to instill visual reasoning skills. Furthermore, prompt engineering for broader visual behaviors is ignored during RL training (Section 5.2, Table 4).
Figure 2: Without actively reinforcing visually grounded behaviors, RL collapses onto shortcuts that maximize immediate rewards at the expense of richer visual reasoning. Standard CoT and Vanilla GRPO (left and center) exhibit visually ungrounded reasoning, relying on vague references to scene elements (shown in yellow), which often results in incorrect answers (marked in red). In contrast, Visually Grounded RL (right) explicitly references object positions, demonstrating precise spatial grounding (shown in blue) and more often producing correct reasoning outcomes (marked in green). See Section 5.3 for further analysis.
Why does standard RL fail here? We hypothesize that the failure mode of RL comes down two interrelated issues: (1) the initial sampling distribution of pretrained VLMs heavily biases the model toward abstract, language-based strategies rather than region-level analysis (as shown in Takeaway 1 above), (2) since standard RL provides rewards based solely on final correctness and general formatting, it amplifies behaviors that attain high rewards irrespective of how they actually maximize reward. Without actively biasing the initial policy and reinforcing visually grounded behaviors, RL naturally collapses onto shortcuts that maximize immediate rewards at the expense of richer visual reasoning. Thus, we hypothesize that encouraging explicit grounding biases the model’s reasoning distributions toward exploring relevant visual regions, setting meaningful visual subgoals, verifying visual hypotheses, and effectively backtracking.
4 Visually Grounded Reinforcement Learning (ViGoRL)
The analysis in Section 3.1 shows that naïve RL on small VLMs may degrade into ungrounded reasoning strategies, and that incentivizing grounding may improve visual reasoning behaviors that improve generalization. To incorporate explicit visual grounding into the reasoning process, we redefine each reasoning step as a tuple: $n_{t}=\langle s_{t},(x_{t},y_{t})\rangle$, where $s_{t}$ is a textual thought and $(x_{t},y_{t})$ anchors it to a specific image location. The full trajectory becomes $\tau=[n_{1},\dots,n_{T},a]$. This modifies the original factorization from Section 3:
$\pi_{\theta}(\tau\mid I,q)=\left(\prod_{t=1}^{T}\pi_{\theta}(n_{t}\mid I,q,n_{<t})\right)\cdot\pi_{\theta}(a\mid I,q,n_{\leq T}),$
where each $n_{t}$ now includes both the reasoning step and its visual grounding. By introducing this grounding constraint, we explicitly guide the model to systematically reference specific image locations as evidence for its reasoning. This grounding incentivizes the model to iteratively explore and verify distinct visual regions, formulate and ground intermediate subgoals visually, and revisit prior regions when uncertainty or errors arise.
4.1 Our Approach for Grounded Reinforcement Learning
Building on our findings in Section 3.1, we introduce a comprehensive approach that directly addresses the ungrounded reasoning patterns in VLMs. We propose a two-stage pipeline to incorporate grounded reasoning as detailed in the previous section: (1) warm-start supervised finetuning that biases the model toward generating structured reasoning chains with explicit spatial grounding, followed by (2) reinforcement learning that systematically refines these grounded behaviors. Finally, we extend our approach to multi-turn RL (Section 4.3.1), enabling fine-grained visual feedback at each reasoning step. This pipeline yields models that invest test-time compute into examining diverse image regions—precisely the behaviors our analysis showed were absent in standard VLM reasoning.
Figure 3: Overview of the ViGoRL approach. (Left) We use MCTS with a teacher model to generate reasoning chains grounded in specific image regions. (Middle) These reasoning trees are linearized and used for supervised fine-tuning (SFT) to train a base model. (Right) We apply GRPO with an outcome-based reward to further refine the grounded reasoning.
4.2 Warm-Start Data Generation via MCTS
MCTS with Visual Grounding. We employ MCTS to generate grounded reasoning traces, where each node is a reasoning step $n_{t}=\langle s_{t},(x_{t},y_{t})\rangle$, anchoring thought $s_{t}$ to image coordinates $(x_{t},y_{t})$ (Figure 3, left panel). At each iteration: (1) Selection: nodes are traversed using UCB, prioritizing high-value, under-explored paths; (2) Expansion: the VLM samples new grounded steps $\langle s_{t},(x_{t},y_{t})\rangle$ for unexplored nodes, each referencing a distinct image region; (3) Simulation: rollouts are performed by recursively generating visually grounded steps until a terminal answer is produced; (4) Backpropagation: a judge scores terminal nodes, with rewards propagated up the path. This process ensures efficient exploration of promising image regions and reasoning steps. We provide additional details on our MCTS procedure in Appendix A5.5.
Why MCTS? Even large models (Qwen2.5-VL-72B) deployed with standard prompting explore only 2–3 regions, set few subgoals, and never backtrack (Table 5). Purely human-curated traces are costly to collect at scale, while linear rollouts cannot enforce the iterative exploration and corrective loops we desire. Empirically, distillation from such linear rollouts without MCTS leads to degraded generalization, performing worse on out-of-distribution spatial tasks after GRPO training (Table 4). In contrast, MCTS lets us systematically search the space of grounded reasoning steps, trading off exploration breadth and depth, and cheaply generate thousands of richly annotated paths that exhibit behaviors like wide exploration, through early branching to cover diverse image regions, and backtracking, by abandoning failing branches and revisiting alternatives.
Moreover, since our reasoning steps are already defined as $(s_{t},(x_{t},y_{t}))$ tuples, they map naturally onto MCTS nodes and transitions, making it both conceptually clean and computationally efficient.
Teacher-guided Search. We employ a frozen, high-capacity teacher (Qwen2.5-VL-72B) to expand each node in the MCTS tree. At node $n$, we prompt the teacher to either (a) generate a new grounded reasoning step $s_{n}$ with coordinate, or (b) emit a candidate answer. We score leaf answers by correctness and backpropagate to guide tree expansion. From 1,500 prompts we derive $\sim$30k high-quality reasoning traces—a dataset orders of magnitude smaller than typical SFT corpora, but densely packed with exploration, verification, and backtracking behaviors.
Linearization for SFT. We then linearize selected root-to-leaf paths into two types of training examples: 1) Direct chains: successful trajectories leading to the correct answer with no detours, and 2) Corrected chains: trajectories where an initial rollout fails, triggers a “wait, that seems off” backtrack, and proceeds to the correct solution (Figure 3, left panel).
We denote the VLM finetuned on these MCTS chains as $\pi_{\theta_{0}}$ (Figure 3, middle). After fine-tuning, we commonly observe reasoning chains with dense visual subgoal setting, visual verification, broad region exploration, and visual backtracking. A representative trace can be seen in Appendix A9.2.
4.3 Reinforcement Learning with Spatially-Grounded Reasoning Steps
Although $\pi_{\theta_{0}}$ imitates high-quality traces, it does not reason optimally for new queries. We therefore apply Group Relative Policy Optimization (GRPO; 50) to directly maximize task reward while preserving fluency and grounding (Figure 3, right). More details can be found in Appendix A5.1.
Reward Design. We wish to incentivize reasoning traces that include explicit grounding coordinates, and reasoning traces that lead to the final answer. Our total reward is a weighted sum: $R(\mathcal{T})=\lambda_{\text{fmt}}r_{\text{fmt}}+\lambda_{\text{task}}r_{\text{task}},$ where $r_{\text{fmt}}$ encourages valid and interpretable output format, and $r_{\text{task}}$ captures task-specific correctness. Importantly, along with checking outputs are formatted with correct <think> and <answer> tags, we award $+1$ $r_{\text{fmt}}$ only if all coordinate references are valid. We provide more details on task reward in Appendix A5.2.
4.3.1 Multi-Turn Reinforcement Learning for Visual Feedback
Grounded reasoning chains generated by our MCTS procedure already push the model to look at many regions, but the visual encoder still processes the same globally resized image at every step. Fine-grained cues (small text, icons, object boundaries) are therefore blurred away, potentially limiting the benefit of additional reasoning if image region details cannot be percieved by the model. Inspired by how humans zoom in after selecting a candidate region, we let the model request a higher-resolution crop, $o_{t}$, after predicting a coordinate. This “interactive microscope” supplies fresh evidence at a detail level impossible to encode in the initial global view.
Multi-turn warm start: from single-turn chains to dialogs. To prepare the model for multi-turn rollouts, we convert single-step MCTS-derived reasoning traces from Section 4.2 into dialog-style traces. Given a linearized MCTS trace $\tau=[(s_{1},\mathbf{p}_{1}),\dots,(s_{T},\mathbf{p}_{T}),a]$, we convert it into a dialog:
  1. 1.
    At turn $t$, the model generates a textual thought $s_{t+1}$ (tagged <think></think>).
  2. 2.
    The model then emits <tool_call> {“name”: “crop”, “arguments”: {“coordinate”: $\mathbf{p}_{t}$}} </tool_call>, or the answer (tagged <answer></answer>), then the round terminates.
  3. 3.
    If a function call, the environment responds with $o_{t}$, with tag <observation> containing a $w\times w$ crop centered at $\mathbf{p}_{t}$, resized to $r\times r$ pixels, followed by </observation> and the loop repeats.
We fine-tune the base model on these multi-turn traces to initialize it for multi-step GRPO with visual feedback. We then apply GRPO, allowing the model to roll out full multi-turn dialogs, with crop feedback upon tool call outputs.
Reward Design for Multi-Turn RL. Multi-turn settings introduce new failure modes: repeating coordinates, skipping tool use, or violating dialog structure. To mitigate these, we define a composite format reward: (1) Grammar reward $r_{\text{grammar}}$: 1 if the dialog obeys a strict tag automaton: <think></think><tool_call></tool_call><observation></observation> → repeat or <answer></answer>. The dialog must end with a complete </answer> and contain no malformed or out-of-order tags. (2) Diversity bonus $r_{\text{div}}$: +0.2 for each sufficiently distinct coordinate in tool calls ($\geq$ 10px from all previous), up to 4 times. Finally, this leads to the overall format reward: $r_{\text{fmt}}=r_{\text{grammar}}+r_{\text{div}}$. Additional multi-turn RL details can be found in Appendix A5.3.
5 Experiments
We evaluate our model, ViGoRL, on spatial reasoning and web grounding tasks, comparing against baseline and ablation variants to understand the contributions of grounding, MCTS-based warm-start supervision, and GRPO. Our results demonstrate significant gains in visual reasoning and web-based tasks when models are trained with our grounded reasoning training recipe. We investigate the following research questions:
[…]
5.1 RQ1: How much does grounded reasoning help on visual reasoning tasks?
Baselines. We compare against the following baselines:
1. Method Comparison: We compare ViGoRL to baselines utilizing the same training data and base model (Qwen-2.5-VL) as used in our method, but differing in their training recipe: 1) SFT-direct: Supervised fine-tuning on our trajectory dataset using final answers only. 2) Vanilla GRPO: GRPO applied to the base model with standard rewards for <think> and <answer> formatting, and answer correctness, similar to previous work [37, 38].
2. General Proprietary and Open-Source Models: General-purpose vision-language models with closed-weights accessible through APIs [1] and open-weight models [35, 7, 8, 63, 3, 59].
3. VLM Tool-Using Pipelines: We compare against prompt-based models that explicitly decompose visual tasks into sequences of subgoals or intermediate steps [67, 19, 23, 88, 76, 66].
4. Web-Grounding Models: Models trained to specialize in web grounding tasks, through large-scale supervised finetuning and instruction tuning on curated web data [6, 21, 70, 34, 15, 45], reinforcement learning of chain of thought with outcome reward [39], or human-in-the-loop action and chain of thought annotations collected from live webpages [49].
5.1.1 Visual Reasoning Evaluations.
Spatial Reasoning. We evaluate on SAT-2 [48] validation (4,000 questions across 5 categories), BLINK [11] (depth ordering, multi-view reasoning, spatial relationships), and ROBOSPATIAL-configuration and compatibility split (228 questions on real-world RGBD scenes) [53]. These benchmarks test generalization to novel environments, objects, and language configurations. Accuracy is measured via multiple-choice answer matching.
GUI Understanding. For grounding evaluation, we use ScreenSpot v2 [70, 6] (single-step localization), and ScreenSpot Pro [31] (high-resolution professional environments). To test small element grounding in low resolution, we additionally evaluate on ScreenSpot Pro with images downsampled to a lower resolution of 1920x1920, which we call ScreenSpot-Pro-LR (results in Appendix Table A6). Performance is measured by checking if predicted coordinates fall within target element bounding boxes. We also test on VisualWebArena [29] live web evaluation, using the visual-only configuration where models receive set-of-marks annotated webpage screenshots without HTML or text inputs. Task success is determined automatically by checking for specific criteria in the agent trajectory (see Koh et al. [29] for details).
Visual Search. The visual search benchmark VBench tests fine-grained visual understanding using 191 high-resolution images from the SA-1B dataset (avg. 2246×1582). It includes 115 attribute recognition samples (e.g., color, material) and 76 spatial relationship samples, evaluating models’ ability to analyze detailed object properties and relative positions in complex scenes.
5.1.2 Results
Table 1: Performance (mean $\pm$ 95% CI) on ScreenSpot and VisualWebArena. Multi-t = multi-turn RL with visual feedback.
Model ScreenSpot ScreenSpot VWA
-V2 -Pro (Vision Only)
Proprietary Models
GPT-4o 18.1 0.8 19.8 (w/ text)
Claude Comp. Use - 17.1
Open-source Models
Qwen2-VL-7B - 1.6 2.9 (w/ text)
Kimi-VL-16B-MoE 92.8 34.5
Web Grounding Models
SeeClick 55.1 1.1
CogAgent-18B - 7.7
OS-Atlas-4B 71.9 3.7
OS-Atlas-7B 84.1 18.9
ShowUI-2B - 7.7
UGround-7B - 16.5
UGround-V1-7B 31.1
UI-TARS-2B 84.7 27.7
UI-R1-3B 17.8
ICAL-7B - 8.2 (w/ text)
Method Comparison
Qwen2.5-VL-3B 68.4% (±2.6) 9.3% (±1.4) 4.2% (±1.3)
+ SFT direct 80.6% (±2.2) 25.0% (±2.1) 4.5% (±1.4)
+ Vanilla GRPO 84.4% (±2.0) 29.0% (±2.2) 4.8% (±1.4)
ViGoRL-3b (Ours) 86.5% (±1.9) 31.1% (±2.3) 6.4% (±1.5)
ViGoRL-3b (Multi-t) 86.1% (±1.9) 32.3% (±2.4) -
Qwen2.5-VL-7B 73.6% (±2.4) 14.8% (±1.8) 5.5% (±1.5)
ViGoRL-7b (Ours) 91.0% (±1.6) 33.1% (±2.3) 11.2% (±2.1)
Table 2: Accuracy (mean ± 95% CI) for spatial reasoning.
Model SAT-2 Val BLINK RoboSpatial
Proprietary Models
GPT-4o 60.0% 76.2%
GPT-4 Turbo 54.6%
Claude 3 Opus 44.1%
Gemini Pro 1.0 45.2%
General Open-source Models
LLaVA v1.6 34B 46.8%
instructBLIP 13B 42.2%
Molmo - 67.1%
Method Comparison
Qwen2.5VL-3B 46.1% (±1.5) 44.4% (±2.3) 54.4% (±6.5)
+ SFT direct 58.3% (±1.5) 46.4% (±2.3) 62.3% (±6.3)
+ Vanilla GRPO 50.0% (±1.6) 46.5% (±2.3) 69.7% (±6.0)
ViGoRL-3b (Ours) 62.9% (±1.5) 48.5% (±2.3) 67.1% (±6.1)
Qwen2.5-VL-7B 52.6% (±1.6) 52.9% (±2.3) 46.7% (±9.5)
ViGoRL-7b (Ours) 67.5% (±1.5) 54.1% (±2.3) 76.4% (±7.5)
Grounded reasoning improves spatial accuracy. As shown in Table 2, on a variety of spatial understanding benchmarks, ViGoRL significantly outperforms baselines without grounded reasoning. On SAT-2 test set, ViGoRL-3B achieves 62.9% accuracy—an improvement of +16.8 points over the base model and +12.9 points over Vanilla GRPO. Similar trends hold on the out of distribution benchmarks of RoboSpatial (67.1%) and BLINK (48.5%), with ViGoRL-3B outperforming SFT Direct by 2.3% and 4.8% on BLINK and RoboSpatial, respectively, and vanilla GRPO by 2.0% on BLINK. At the 7B scale, ViGoRL-7B further improves SAT-2 accuracy to 67.5%, demonstrating the method’s scalability and effectiveness across model sizes. We observe the same trend in BLINK and RoboSpatial, with improved accuracies of 54.1% and 76.4%, respectively. Importantly, these gains come without sacrificing general visual-language capabilities, as our method maintains performance on standard out-of-distribution VLM benchmarks (Table A7).
Grounded reasoning helps localize complex web elements. We evaluate performance on vision-only web interfaces (Table 2), where models must resolve ambiguous UI instructions via grounded image understanding. ViGoRL-3B outperforms both the base model, direct answer SFT, and vanilla GRPO across all tasks. For instance, on ScreenSpot-Pro, requiring grounding of small elements in high resolution webpage images, accuracy improves 21.8 points over the base model, 6.1 points over SFT direct, and 2.1 points over vanilla GRPO. ViGoRL-7B achieves 91.0% on ScreenSpot-V2 and 33.1% on ScreenSpot-Pro, outperforming open-source VLMs of comparable size, including OS-Atlas and UGround variants, which finetune on order of magnitudes more GUI grounding examples (up to 13M training samples).
Table 3: Accuracy (mean ± 95% CI) for visual search on VBench.
Model Name VBench
Proprietary Models
Gemini-Pro 48.2
GPT-4V 55.0
GPT-4o 66.0
VLM Tool-Using Pipelines
VisProg 41.4
VisualChatGPT 37.6
MM-React 41.4
Sketchpad-GPT-4o 80.3
IVM-Enhanced GPT-4V 81.2
Open-source Models
LLaVA-1.5-7B 48.7
LLaVA-1.6-13B 61.8
SEAL 74.8
Qwen2.5-7B-VL 78.0 (±5.9)
Multi-t ViGoRL-7B (Ours) 86.4 (±5.0)
Qwen2.5-3B-VL 74.2 (±6.2)
Multi-t ViGoRL-3B (Ours) 81.2 (±5.6)
ViGoRL improves accuracy on live visual web evaluation. On VisualWebArena, which requires live interaction with web pages using only set-of-marks image inputs (no access to HTML or underlying text), ViGoRL outperforms direct SFT and vanilla GRPO. Despite relying solely on images, ViGoRL surpasses the previous state-of-the-art for the same model size, ICAL [49], by 3.0%—even though ICAL has access to textual set-of-marks inputs derived from HTML.
Multi-turn RL with visual feedback improves visual search and small element detection. On VBench, ViGoRL-7B significantly outperforms both proprietary models and open-source tool-using pipelines. As shown in Table 3, our method reaches 86.4%, surpassing proprietary VLMs like GPT-4o (66.0%) and Gemini-Pro (48.2%), and even advanced tool-using pipelines like VisProg (41.4%) Sketchpad-GPT-4o (80.3%).
Our model with multi-turn RL that incorporates zoomed-in visual feedback additionally shows significant improvements in small element grounding tasks. It achieves 32.3% accuracy on ScreenSpot-Pro, a 1.2 percentage point improvement over our best non-visual feedback model variant (31.1%) (Table 2). In low-resolution environments, the zooming capability delivers even more substantial relative gains, outperforming our best non-visual feedback method by 2.4% (Table A6). These gains illustrate that our approach enables more robust and precise visual reasoning through structured visual grounding and dynamic visual feedback interactions.
5.2 RQ2: Ablation Studies
Table 4: Ablation results on SAT-2 Val and BLINK. Top-1 accuracy (±95% CI) with relative change. Model never produced correct formatting. Gnd = Explicit grounding in the reasoning steps. SFT = SFT Direct pretraining. Distill = warm-start teacher distillation.
Variant GRPO MCTS Gnd Distill SAT-2 BLINK
Full (ours) 62.93 (±1.50) 48.50 (±2.33)
– GRPO 58.83 (±1.53) (–4.10%) 44.97 (±2.32) (–3.53%)
– MCTS N/A N/A
– Grounded 58.69 (±1.53) (–4.28%) 45.44 (±2.32) (–3.06%)
Distill w/o MCTS 63.28 (±1.49) (+0.35%) 46.18 (±2.32) (–2.32%)
+ SFT direct pre-train 63.26 (±1.49) (+0.33%) 48.56 (±2.33) (+0.06%)
To understand which components drive performance, we conduct targeted ablations (Table 4):
1. GRPO remains important. Removing GRPO reduces performance by 4.1 points on SAT-2 and 3.5 points on BLINK, confirming the importance of RL refinement.
2. MCTS-generated warm-start is essential. Without structured traces from MCTS, the model fails to emit valid outputs, underscoring the need for scaffolded learning signals.
3. Explicit grounding helps. Running our same method without introducing explicit grounding in the warm start data resulted in a 3.4 point drop on BLINK, even with the same MCTS warm start and GRPO recipe.
4. Teacher distillation without MCTS preserves in-distribution performance but degrades out-of-distribution generalization. When using warm-start data from successful teacher linear rollouts (which contain grounded reasoning steps but no search), SAT-2 validation accuracy remains nearly unchanged (63.26% vs. 62.93%). However, performance on the more challenging out-of-distribution BLINK benchmark drops significantly by 2.32%.
5. SFT direct pretraining adds little. Adding an additional SFT stage to directly predict the answer before warm start training and GRPO yields marginal gains (<1 point), reinforcing the idea that grounded thought processes drive improvement.
5.3 RQ3: What visual reasoning behaviors are amplified by grounded reasoning?
Table 5: Average visual behaviors per example.
Regions Grounded Visually
Model Explored Subgoals Verify Backtrack Acc.
Qwen2.5-VL-72B (Zero-Shot)
Standard CoT 2.3 0.27 0.27 0.00 0.65
Qwen2.5-VL-3B (Zero-Shot)
Standard CoT 1.44 0.07 0.14 0.00 0.38
Qwen2.5-VL-3B (RL-tuned)
Vanilla GRPO 1.8 0.00 0.17 0.00 0.48
Ours 3.5 1.1 0.39 0.47 0.64
w/o grounding 1.7 0.02 0.27 0.02 0.58
Following previous work on behavioral coding in language models [13], we code VLM behavior when tasked with the SAT-2 spatial reasoning benchmark, which requires synthesizing evidence across multiple image regions, in two scenarios: (1) Zero-shot prompting with "think step-by-step" instructions without explicit grounding, and (2) RL-tuned models including vanillo GRPO, ViGoRL, and an ablated version without explicit grounding incentives. We quantify reasoning behaviors on 300 representative SAT-2 samples using GPT-4o evaluation, measuring visual regions explored, subgoal setting, verification, and backtracking. Additional study details can be found in Appendix A6.
Findings. As shown in Table 5, explicit grounding substantially amplifies visual reasoning behaviors. ViGoRL explores more visual regions (3.5 vs. 1.44/1.8 in zero-shot/vanilla GRPO) and demonstrates dramatic increases in grounded subgoals (1.1 vs. 0.07, 15× higher) and visual verification (0.39 vs. 0.14, 3× higher). It uniquely develops visual backtracking behavior (0.47) absent in all baselines. These enhanced behaviors enable our 3B parameter model to achieve accuracy (0.64) comparable to the 72B model (0.65). The ablation study confirms these benefits stem specifically from explicit grounding incentives, as removing them causes substantial regression in all measured behaviors, particularly regions explored (1.7 vs. 3.5) and grounded subgoals (0.02 vs. 1.1).
5.4 RQ4: Human Evaluation Shows Accuracy and Helpfulness of Grounded Reasoning
Figure 4: Human evaluation of grounded reasoning. Participants judged the grounded predictions as both accurate and helpful when correct.
To assess our model’s grounded reasoning traces, we conducted a human study evaluating whether predicted coordinates (1) correctly referred to the intended image region, and (2) helped participants understand the associated reasoning step (details are shown in Appendix A7).
Findings. As shown in Figure 4, 72.8% of predictions were judged as accurately referring to the described region (95% CI: [66.8, 78.7], N=20). On a 5-point Likert scale, participants rated the helpfulness of the highlighted region at 3.35 on average (95% CI: [3.03, 3.68], N=10). Helpfulness increased substantially when the prediction was correct (3.81; 95% CI: [3.51, 4.10]) and dropped when incorrect (2.26; 95% CI: [1.57, 2.94]).
These findings indicate that accurate spatial grounding meaningfully improves human interpretability and usefulness of the model’s reasoning process, indicating that improvements in accuracy of reasoning step grounding can also improve human interpretability.
6 Conclusion and Discussion
Why does visual grounding help? Our findings suggest that spatially anchoring each reasoning step forces the model to engage in a more structured, human-aligned form of cognition. ViGoRL learns to iteratively reference, inspect, and verify content in specific visual regions – amplifying cognitive behaviors such as subgoal formulation, visual verification, and backtracking.
This model architecture mirrors insights from cognitive science: humans rely on spatial attention and visual routines to decompose complex problems into manageable, perceptually grounded steps [62, 4]. Grounding serves not merely to reduce computational load (as human spatial attention is often characterized), but to scaffold reasoning with external visual structure – effectively using the content of the world as part of the thinking process [78, 20]. We observe similar benefits in models: spatial grounding enables better generalization, especially in out-of-distribution settings, and improves interpretability by making intermediate steps physically referable.
Rather than treating grounding as a visualization tool or auxiliary supervision signal, our results argue for it as a central architectural and algorithmic principle. By training models to reason with deictic reference – pointing, zooming, verifying – future systems may better reflect the iterative, grounded strategies that underlie human problem-solving. This opens up promising directions for building agents that not only reason effectively, but in ways that are queryable, adaptable, and aligned with perceptual experience.
Acknowledgments. This material is based upon work supported by National Science Foundation grants GRF DGE1745016 & DGE2140739 (GS), ONR award N00014-23-1-2415, AFOSR Grant FA9550-23-1-0257, and DARPA No. HR00112490375 from the U.S. DARPA Friction for Accountability in Conversational Transactions (FACT) program. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the United States Army, the National Science Foundation, or the United States Air Force.
This research project has benefitted from the Microsoft Accelerate Foundation Models Research (AFMR) grant program through which leading foundation models hosted by Microsoft Azure along with access to Azure credits were provided to conduct the research.

### v2

@@
 This hypothesis is inspired by the fact that humans systematically shift their restricted gaze to selectively gather and integrate task-relevant information when reasoning about the world [77, 24, 76].
-Grounding may serve a similar role in models, functioning as a spatial attention mechanism that enables accurate feature binding [61] and supports deictic reference [4] to simplify multi-step reasoning through localized perceptual anchoring.
+Grounding may serve a similar role in models, functioning as a spatial attention mechanism that enables accurate feature binding [63, 5, 6] and supports deictic reference [4] to simplify multi-step reasoning through localized perceptual anchoring.
 We move beyond prompt-based reasoning, proposing that learning to compose reasoning steps explicitly anchored in image coordinates induces structured region-level behaviors that support improved generalization in visual tasks.
@@
 where the reward $R(\tau)$ typically includes correctness of the final answer, and proper adherence to structured reasoning formats.
-3.1 Do Current RL Recipes Amplify VLM Behaviors That Support Visual Performance?
+3.1 Do Current RL Recipes Amplify VLM Behaviors That Support Visual Reasoning?
 It has been shown that RL on chains of thought alone cannot necessarily induce new behaviors from scratch; it can only amplify or chain reasoning primitives that are already present in the base model’s sampling distribution [15, 84].
@@
 4.3 Reinforcement Learning with Spatially-Grounded Reasoning Steps
-Although $\pi_{\theta_{0}}$ imitates high-quality traces, it does not reason optimally for new queries.
-We therefore apply Group Relative Policy Optimization (GRPO; 50) to directly maximize task reward while preserving fluency and grounding (Figure 3, right).
+Although $\pi_{\theta_{0}}$ imitates high‑quality traces, it does not reason optimally for new queries.
+We therefore apply Group Relative Policy Optimization (GRPO; 52) to directly maximize task reward while preserving fluency and grounding (Figure 3, right).
 More details can be found in Appendix A5.1.
@@
 Our total reward is a weighted sum: $R(\mathcal{T})=\lambda_{\text{fmt}}r_{\text{fmt}}+\lambda_{\text{task}}r_{\text{task}},$ where $r_{\text{fmt}}$ encourages valid and interpretable output format, and $r_{\text{task}}$ captures task-specific correctness.
-Importantly, along with checking outputs are formatted with correct <think> and <answer> tags, we award $+1$ $r_{\text{fmt}}$ only if all coordinate references are valid.
+Importantly, along with checking outputs are formatted with correct <think> and <answer> tags, we award +1+1 $r_{\text{fmt}}$ only if all coordinate references are valid.
 We provide more details on task reward in Appendix A5.2.
@@
 Grounded reasoning chains generated by our MCTS procedure already push the model to look at many regions, but the visual encoder still processes the same globally resized image at every step.
-Fine-grained cues (small text, icons, object boundaries) are therefore blurred away, potentially limiting the benefit of additional reasoning if image region details cannot be percieved by the model.
+Fine-grained cues (small text, icons, object boundaries) are therefore blurred away, potentially limiting the benefit of additional reasoning if image region details cannot be perceived by the model.
 Inspired by how humans zoom in after selecting a candidate region, we let the model request a higher-resolution crop, $o_{t}$, after predicting a coordinate.
@@
 2) Vanilla GRPO: GRPO applied to the base model with standard rewards for <think> and <answer> formatting, and answer correctness, similar to previous work [39, 40].
+Our model and all baselines decode with a temperature of 0.5 during evaluation.
 2. General Proprietary and Open-Source Models: General-purpose vision-language models with closed-weights accessible through APIs [1] and open-weight models [37, 9, 10, 65, 3, 61].
@@
 Accuracy is measured via multiple-choice answer matching.
+We use a max side length of 1260 pixels while keeping aspect ratio for evaluation.
 GUI Understanding.
@@
 [31] for details).
+We use a resize image to have maximum pixel size of 3600x3600.
 Visual Search.
@@
 It includes 115 attribute recognition samples (e.g., color, material) and 76 spatial relationship samples, evaluating models’ ability to analyze detailed object properties and relative positions in complex scenes.
+We use a resize image to have maximum pixel size of 3600x3600.
 5.1.2 Results
@@
 Method Comparison
-Qwen2.5-VL-3B 68.4% (±2.6) 9.3% (±1.4) 4.2% (±1.3)
-+ SFT direct 80.6% (±2.2) 25.0% (±2.1) 4.5% (±1.4)
-+ Vanilla GRPO 84.4% (±2.0) 29.0% (±2.2) 4.8% (±1.4)
-ViGoRL-3b (Ours) 86.5% (±1.9) 31.1% (±2.3) 6.4% (±1.5)
-ViGoRL-3b (Multi-t) 86.1% (±1.9) 32.3% (±2.4) -
-Qwen2.5-VL-7B 73.6% (±2.4) 14.8% (±1.8) 5.5% (±1.5)
-ViGoRL-7b (Ours) 91.0% (±1.6) 33.1% (±2.3) 11.2% (±2.1)
+Qwen2.5-VL-3B 68.4 (±2.6) 23.9 4.2 (±1.3)
++ SFT direct 80.6 (±2.2) 25.0 (±2.1) 4.5 (±1.4)
++ Vanilla GRPO 84.4 (±2.0) 29.0 (±2.2) 4.8 (±1.4)
+ViGoRL-3b (Ours) 86.5 (±1.9) 31.1 (±2.3) 6.4 (±1.5)
+Multi-t ViGoRL-3b (Ours) 86.1 (±1.9) 32.3 (±2.4) -
+Qwen2.5-VL-7B 73.6 (±2.4) 29.0 5.5 (±1.5)
+ViGoRL-7b (Ours) 91.0 (±1.6) 33.1 (±2.3) 11.2 (±2.1)
 Table 2: Accuracy (mean ± 95% CI) for spatial reasoning.
-Model SAT-2 Val BLINK RoboSpatial
+Model SAT‑2 Val BLINK RoboSpatial
 Proprietary Models
-GPT-4o 60.0% 76.2%
-GPT-4 Turbo 54.6%
-Claude 3 Opus 44.1%
-Gemini Pro 1.0 45.2%
+GPT-4o 60.0 76.2
+GPT-4 Turbo 54.6
+Claude 3 Opus 44.1
+Gemini Pro 1.0 45.2
 General Open-source Models
-LLaVA v1.6 34B 46.8%
-instructBLIP 13B 42.2%
-Molmo - 67.1%
+LLaVA v1.6 34B 46.8
+instructBLIP 13B 42.2
+Molmo - 67.1
 Method Comparison
-Qwen2.5VL-3B 46.1% (±1.5) 44.4% (±2.3) 54.4% (±6.5)
-+ SFT direct 58.3% (±1.5) 46.4% (±2.3) 62.3% (±6.3)
-+ Vanilla GRPO 50.0% (±1.6) 46.5% (±2.3) 69.7% (±6.0)
-ViGoRL-3b (Ours) 62.9% (±1.5) 48.5% (±2.3) 67.1% (±6.1)
-Qwen2.5-VL-7B 52.6% (±1.6) 52.9% (±2.3) 46.7% (±9.5)
-ViGoRL-7b (Ours) 67.5% (±1.5) 54.1% (±2.3) 76.4% (±7.5)
+Qwen2.5VL-3B 46.1 (±1.5) 44.4 (±2.3) 54.4 (±6.5)
++ SFT direct 58.3 (±1.5) 46.4 (±2.3) 62.3 (±6.3)
++ Vanilla GRPO 50.0 (±1.6) 46.5 (±2.3) 69.7 (±6.0)
+ViGoRL-3b (Ours) 62.9 (±1.5) 48.5 (±2.3) 67.1 (±6.1)
+Qwen2.5-VL-7B 52.6 (±1.6) 52.9 (±2.3) 46.7 (±9.5)
+ViGoRL-7b (Ours) 67.5 (±1.5) 54.1 (±2.3) 76.4 (±7.5)
 Grounded reasoning improves spatial accuracy.
@@
 VisualChatGPT 37.6
-MM-React 41.4
+$M$-React 41.4
 Sketchpad-GPT-4o 80.3
@@
 SEAL 74.8
-Qwen2.5-7B-VL 78.0 (±5.9)
-Multi-t ViGoRL-7B (Ours) 86.4 (±5.0)
-Qwen2.5-3B-VL 74.2 (±6.2)
-Multi-t ViGoRL-3B (Ours) 81.2 (±5.6)
+Qwen2.5-7B-VL 78.0 (±3.0)
+Multi-t ViGoRL-7B 86.4 (±2.7)
+Qwen2.5-3B-VL 74.2 (±3.1)
+ViGoRL-3B (Ours) 79.1 (±3.0)
+Multi-t ViGoRL-3B 81.2 (±2.8)
+w/o diversity reward 78.0 (±3.0)
+w/ bounding box outputs 81.2 (±2.8)
 ViGoRL improves accuracy on live visual web evaluation.
@@
 Adding an additional SFT stage to directly predict the answer before warm start training and GRPO yields marginal gains (<1 point), reinforcing the idea that grounded thought processes drive improvement.
+6.
+Pointing is sufficient for visual search.
+As shown in Table 3, we observe no difference in V*Bench accuracy when switching to training ViGoRL to output variable size bounding box crops as opposed to the fixed-size point cropping used in ViGoRL.
+Fixed crops usually provide enough context and resolution for most cases, and ViGoRL can handle larger regions through multiple overlapping crops that together capture necessary information.
+Adaptive cropping could be more efficient but is limited by imprecise bounding boxes and inconsistent aspect ratios.
 5.3 RQ3: What visual reasoning behaviors are amplified by grounded reasoning?